Evaluate:


sin 30⁰ cos 45⁰ − cos 30 ⁰ sin 45⁰


1
Expert's answer
2022-03-29T11:29:43-0400

sin 30⁰ cos 45⁰ − cos 30 ⁰ sin 45⁰"=\frac{1}{2}*\frac{\sqrt{2}}{2}-\frac{\sqrt{3}}{2}*\frac{\sqrt{2}}{2}=-\frac{\sqrt{2}(\sqrt{3}-1)}{4}."